OpenStudy (anonymous):

find the length of the curve: y = ln(cos x) 0

OpenStudy (anonymous):

\[s = \int\limits_0^{\frac{\pi}{3}} \sqrt{1+ |f'(x)|^2}dx = \int\limits_0^{\frac{\pi}{3}} \sqrt{1+ \tan(x)^2}dx = \log(2+\sqrt{3}) = 1.31\]

OpenStudy (anonymous):

how'd you get log(2 + root3?

OpenStudy (jamesj):

Right. Now, it's not obvious how you did the integration, although it is helps a lot to remember that 1 + tan^2 x = sec^2 x. So you need the integral of sec x

OpenStudy (anonymous):

right now i have the integral of (secx^2)^1/2

OpenStudy (jamesj):

Which is sec x

OpenStudy (anonymous):

the integral of sec?

OpenStudy (jamesj):

\[\int\limits \sec x \ dx = \int\limits \frac{\sec x (\sec x + \tan x)}{\sec x + \tan x} \ dx\]

OpenStudy (anonymous):

but its secx^2 not just sec x

OpenStudy (jamesj):

No it's not. It is sec x, regardless of whatever notation cptnc used.

OpenStudy (anonymous):

The sqrt cancels the ^2 out

OpenStudy (anonymous):

oh duh!

OpenStudy (anonymous):

okay so the answer is ln(2 + root3)